# Learning Design-Score Manifold to Guide Diffusion Models for Offline Optimization

### v1

@@ Methods
In this section, we begin by defining the setting of offline optimization, encompassing both offline SOO and offline MOO. Subsequently, we delve into the core components of our approach: (i) training the diffusion model to learn the manifold and (ii) guidance generation with the preferred condition, and then introduce the baseline methods.
Problem Definition
Offline SOO aims to identify the optimal design $\bm{x}^{*}=\arg\min_{\bm{x}\in\mathcal{X}}f(\bm{x})$ using only a pre-collected offline dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$, where $\bm{x}_{i}$ denotes a specific design (also referred to as a solution) and $y_{i}=f(\bm{x}_{i})$ represents its corresponding score (or objective value). Offline MOO aims to identify a set of designs that achieve optimal trade-offs among conflicting objectives using a pre-collected dataset $\mathcal{D}=\left\{(\bm{x}_{i},\bm{y}_{i})\right\}_{i=1}^{N}$, where $\bm{y}_{i}$ denotes the vector of scores corresponding to design $\bm{x}_{i}$. The problem is defined as [29]: Find $\text{Find }\bm{x}^{*}\in\mathcal{X}\text{ such that }\nexists\bm{x}\in\mathcal{X}\text{ with }\bm{f}(\bm{x})\prec\bm{f}(\bm{x}^{*}),$ where $\prec$ denotes Pareto dominance. A design $\bm{x}^{\prime}$ is said to Pareto dominate another design $\bm{x}$, denoted as $\bm{f}(\bm{x}^{\prime})\prec\bm{f}(\bm{x})$, if $\exists i\in\{1,\ldots,m\},f_{i}(\bm{x}^{\prime})<f_{i}(\bm{x})$ and $\forall j\in\{1,\ldots,m\},f_{j}(\bm{x}^{\prime})\leq f_{j}(\bm{x}).$ Namely, $\bm{x}^{\prime}$ is superior to $\bm{x}$ in at least one objective while being at least as good in all others. A design $\bm{x}^{*}$ is Pareto optimal if no other design $\bm{x}\in\mathcal{X}$ Pareto dominates $\bm{x}^{*}$. The set of all Pareto optimal designs is referred to as the Pareto set (PS), and the set of their scores $\{\bm{f}(\bm{x}^{*})\mid\bm{x}^{*}\in\text{PS}\}$ constitutes the Pareto front. The goal of offline MOO is to identify the PS using a pre-collected dataset, thereby achieving optimal trade-offs among conflicting objectives.
Learning Design-Score Manifold via Diffusion Models
Diffusion models are deep generative models that learn to reverse a gradual noising process, transforming random noise into realistic data through iterative denoising. Let $\bm{x}_{t}$ denote the state of a data sample $\bm{x}_{0}$ at time $t\in[0,T]$, where $\bm{x}_{0}$ is drawn from an unknown data distribution $p_{0}(\bm{x})$. Here, $\bm{x}_{t}$ represents a noisy version of $\bm{x}_{0}$ at time $t$, and $\bm{x}_{T}$ corresponds to a point sampled from a prior noise distribution $p_{T}(\bm{x})$, typically chosen as the standard normal distribution $p_{T}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$. The forward diffusion process, also known as the noise-adding process, can be modeled as a stochastic differential equation (SDE) [30]: $d\bm{x}=\mathbf{f}(\bm{x},t)dt+g(t)d\bm{w},$ where $\bm{w}$ denotes the standard Wiener process, $\mathbf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, and $g(t):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient of $\bm{x}_{t}$. The denoising process is defined by the reverse-time SDE: $d\bm{x}=\left[\mathbf{f}(\bm{x},t)-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]dt+g(t)d\tilde{\bm{w}},$ where $dt$ represents an infinitesimal step backward in time, and $d\tilde{\bm{w}}$ is the reverse-time Wiener process.
@@
Our method aims to capture the prior probability distribution of the design-score manifold using a diffusion model, a critical step for guiding the model to solve offline optimization problems. To achieve this, we train an unconditional diffusion model based on a variance-preserving (VP) SDE [30]. The model is trained on a joint design-score dataset $\hat{\mathcal{D}}=\{\hat{\bm{x}_{i}}:\hat{\bm{x}_{i}}=(\bm{x}_{i},\bm{y}_{i})\in\mathbb{R}^{(d+m)}\}_{i=1}^{N}$, where $\bm{x}_{i}$ denotes the design and $\bm{y}_{i}$ represents its corresponding score vector. Specifically, $\hat{\mathcal{D}}$ is constructed by augmenting the original design dataset $\mathcal{D}_{x}=\{\bm{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$ with score information.
@@
where $d\tilde{\bm{w}}$ denotes the reverse-time Wiener process. Using this process, the pre-trained diffusion model can transform a prior noise point $\hat{\bm{x}}_{T}\sim p_{T}(\hat{\bm{x}})$ into a design-score-joint point $\hat{\bm{x}}_{0}=(\bm{x}_{0},\bm{y}_{0})$ on the design-score manifold.
@@
The score function $\nabla_{\hat{\bm{x}}}\log p_{t}(\hat{\mathbf{x}})$ in Equation (2) is approximated by a time-dependent neural network $\bm{s}_{\bm{\theta}}\left(\hat{\mathbf{x}}_{t},t\right)$ with parameters $\bm{\theta}$ based on VP-SDE  [30]. The network is optimized by minimizing the following loss function:
@@
1:  Input: Offline dataset $\mathcal{D}$, scaling threshold $\tau$, preferred scores $\bm{y}_{\text{p}}$, preferred design constraints $[\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$, conditional sample number $K$, unconditional sample number $M$, duplication size $J$, importance ratio $\alpha_{\rm I}$
@@
4:  Augment $\mathcal{D}$ and get $\hat{\mathcal{D}}\leftarrow\left\{\hat{\bm{x}}_{i}:\hat{\bm{x}}_{i}=(\bm{x}_{i},\bm{y}_{i})\right\}_{i=1}^{N}$
@@
11:  for $t=T-1$ to 00 do
12:     Guide to generate $\bm{X}\leftarrow\{\hat{\bm{x}}_{t}\}_{i=1}^{K}$ with E.q.(5) on $\bm{y}_{\text{p}}$ and $[\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$
13:     if $\mathcal{F}(\bm{s}_{\bm{\theta}})>\tau$ then
14:        Duplicate $J$ times on $\hat{\bm{x}}^{(i)}_{t}$ and get $\{\hat{\bm{x}}^{(i,j)}_{t}\}_{j=1}^{J}$ for $i\in[K]$
15:        Compute self-supervised reward $\{\mathcal{R}(\hat{\bm{x}}^{(i,j)}_{t})\}_{j=1}^{J}$ with Eqs. (8,9) for $i\in[K]$
16:        Compute importance $w^{(i,j)}_{t}=\exp(\mathcal{R}(\hat{\bm{x}}^{(i,j)}_{t})/\alpha_{\rm I})$
17:        Resample $j_{i}^{*}=\text{Cat}\left(\{w^{(i,j)}_{t}/\sum_{j=1}^{J}w^{(i,j)}_{t}\}_{j=1}^{J}\right)$ for $i\in[K]$
18:        Update $\hat{\bm{x}}^{(i)}_{t}\leftarrow\hat{\bm{x}}^{(i,j_{i}^{*})}_{t}$ for $i\in[K]$
19:     end if
20:  end for
Bidirectional Guidance Generation on Design-Score Samples
@@
Derivative-Free Guidance via Preferred Scores and Preferred Design
In terms of score-to-design generation, the proposed method enables guided generation of preferred design-score samples through two mechanisms: (1) leveraging preferred scores $\bm{y}_{\text{p}}$ to guide the score component of the generated samples, ensuring alignment with desired performance; and (2) specifying a design constraint range $[\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$ to guide the design component, ensuring design feasibility. This dual-guidance strategy generates samples that satisfy both design and score constraints without requiring an additional surrogate score model.
Concretely, during the reverse process, the guidance on the score component of the generated sample $\hat{\bm{x}}_{t}=(\bm{x}_{t},\bm{y}_{t})$ is formulated as the gradient of the MSE between $\bm{y}_{\text{p}}$ and the current score $\bm{y}_{t}$, i.e., $\nabla_{\bm{y}_{t}}\|\bm{y}_{\text{p}}-\bm{y}_{t}\|_{2}^{2}$. Similarly, the guidance on the design component is formulated as the gradient of the MSE from the current design $\bm{x}_{t}$ to the constraint range $[\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$, i.e., $\nabla_{\bm{x}_{t}}\|\bm{x}_{t}-\text{clip}(\bm{x}_{t},\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}})\|_{2}^{2}$, where $\text{clip}(\bm{x}_{t},\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}})$ ensures that $\bm{x}_{t}$ is projected onto the constraint range.
@@
where $\mathcal{F}(\bm{s}_{\bm{\theta}})\in(0,1]$, $j_{i}^{*}=\arg\min_{j}\|{\bm{x}}_{\text{uc}}^{(i)}-\bm{x}_{\text{train}}^{(j)}\|_{2}$ denotes the index of the nearest training data to ${\bm{x}}_{\text{uc}}^{(i)}$ in $\mathcal{D}=\{\hat{\bm{x}}^{(j)}_{\text{train}}=(\bm{x}^{(j)}_{\text{train}},\bm{y}^{(j)}_{\text{train}})\}_{j=1}^{N}$. A higher value of $\mathcal{F}(\bm{s}_{\bm{\theta}})$ indicates a stronger alignment between the generated and ground-truth scores. Our experiments demonstrate that this metric exhibits a strong correlation with inference-time scaling performance based on the self-supervised reward mechanism. Based on this observation, we implement a conditional logic with $\mathcal{F}(\bm{s}_{\bm{\theta}})>\tau$ to activate the inference-time scaling scheme during guidance generation.
In this way, we establish a general inference-time scaling framework for offline optimization by substituting external reward models in existing schemes with self-supervised rewards. This framework is compatible with various methods, including the IS-based scaling [34] and the FKS-based scaling [35] methods. For example, we illustrate the complete process of ManGO based on IS with the self-supervised reward in Algorithm 1. In our experiments, we specifically assess the integration of the IS and FKS approaches with self-supervised rewards, demonstrating their effectiveness in the context of our framework.
Training and Inference Details
@@
Baseline Methods
We consider existing baseline methods for offline SOO based on three methodological paradigms: (i) Surrogate-based methods: optimizing with surrogate models, including BO-$q$EI [37, 38], CMA-ES [10], REINFORCE [39], Gradient Ascent and its variants of mean ensemble and min ensemble. (ii) Forward-modeling methods: employing advanced neural networks like generative models as surrogate models and integrating with surrogate-based methods, including COMs [16], RoMA [40], IOM [41], BDI [11], ICT [42], Tri-Mentoring [43], PGS [44], FGM [45], Match-OPT [46], and RaM [17]. (iii) Inverse-modeling methods: applying score as a condition to reverse design with generative models, including CbAS [47], MINs [48], DDOM [12], BONET [49], and GTG [50].
All results are normalized using task-specific references for comparison: For Design-Bench with maximization tasks (Table 1), we use standardization normalization for the Superconduct task and min-max normalization for other tasks based on the unobserved dataset’s highest score. For the Off-MOO-Bench with minimization tasks (Tables 2 ), we use min-max normalization with the best HV and IGD values of training datasets.
@@ Acknowledgements
The ManGO’s main idea and training part were done during Tailin’s internship at Huawei Noah’s Ark Lab, and the inference part was done at HKUST. This work was supported in part by the Hong Kong Research Grants Council under the Areas of Excellence scheme grant AoE///E-601///22-R, in part by NSFC///RGC Collaborative Research Scheme grant CRS__\__HKUST603///22, in part by Guangzhou Municipal Science and Technology Project under Grant 2023A03J0011, Guangdong Provincial Key Laboratory of Integrated Communications, Sensing and Computation for Ubiquitous Internet of Things, and National Foreign Expert Project, Project Number G2022030026L. Some icons in Figure 1 are from the website https://www.cleanpng.com/.

### v2

@@ Methods
In this section, we delve into the core components of our approach: (i) training the diffusion model to learn the design-score manifold and (ii) bidirectional guidance generation with the preferred condition. Finally, we present the training and inference settings of our proposed approach.
@@
To better explore unknown design-score pairs, our method aims to capture the prior probability distribution of the design-score manifold using a diffusion model. This is a critical step to improve OOG limitation of backward methods to solve offline optimization problems. To achieve this, we train an unconditional diffusion model based on a variance-preserving (VP) SDE [18]. The model is trained on a joint design-score dataset $\hat{\mathcal{D}}=\{\hat{\bm{x}_{i}}:\hat{\bm{x}_{i}}=(\bm{x}_{i},\bm{y}_{i})\in\mathbb{R}^{(d+m)}\}_{i=1}^{N}$, where $\bm{x}_{i}$ denotes the design and $\bm{y}_{i}$ represents its corresponding score vector. Specifically, $\hat{\mathcal{D}}$ is constructed by augmenting the original design dataset $\mathcal{D}_{x}=\{\bm{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$ with score information.
@@
[…]
@@
$\mathcal{L}(\bm{\theta})=\underset{t}{\mathbb{E}}\left[\lambda(t)\underset{\hat{\bm{x}}_{0}}{\mathbb{E}}\left[\bm{w}(\hat{\bm{x}}_{0})\underset{\hat{\bm{x}}_{t}\mid\hat{\bm{x}}_{0}}{\mathbb{E}}\left[\left\|\bm{s}_{\bm{\theta}}\left(\hat{\bm{x}}_{t},t\right)-\nabla_{\hat{\bm{x}}}\log p_{t}\left(\hat{\bm{x}}_{t}|\hat{\bm{x}}_{0}\right)\right\|_{2}^{2}\right]\right]\right],$ (4)
@@
1:Input: Offline dataset $\mathcal{D}$, scaling threshold $\tau$, preferred scores $\bm{y}_{\text{p}}$, preferred design constraints $\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$, conditional sample number $K$, unconditional sample number $M$, duplication size $J$, importance ratio $\alpha_{\rm I}$
@@
[…]
@@
11:for t=T1t=T-1 to 0 do
12:  if $\mathcal{F}(\bm{s}_{\bm{\theta}})>\tau$ then
13:   Duplicate $J$ times on $\hat{\bm{x}}^{(i)}_{t}$ and get $\bm{X}_{i}=\{\hat{\bm{x}}^{(i,j)}_{t}\}_{j=1}^{J}$ for $i\in[K]$
14:   Guide to reverse $\bm{X}_{i}\leftarrow\{\hat{\bm{x}}_{t}^{i,j}\}_{j=1}^{J}$ with Eq.(5) on $\bm{y}_{\text{p}}$ and $\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$ for $i\in[K]$
15:   Estimate the denoised state $\hat{\bm{x}}^{(i,j)}_{0|t}$ with Eq.(8) for $i,j\in[K],[J]$
16:   Compute self-supervised reward $\mathcal{R}(\hat{\bm{x}}^{(i,j)}_{t})$ with Eq.(9) for $i,j\in[K],[J]$
17:   Compute importance $w^{(i,j)}_{t}=\exp(\mathcal{R}(\hat{\bm{x}}^{(i,j)}_{t})/\alpha_{\rm I})$ for $i,j\in[K],[J]$
18:   Resample $j_{i}^{*}=\text{Cat}\left(\{w^{(i,j)}_{t}/\sum_{j=1}^{J}w^{(i,j)}_{t}\}_{j=1}^{J}\right)$ for $i\in[K]$
19:   Update $\hat{\bm{x}}^{(i)}_{t}\leftarrow\hat{\bm{x}}^{(i,j_{i}^{*})}_{t}$ for $i\in[K]$ and get $\bm{X}\leftarrow\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{K}$
20:  else
21:   Guide to reverse $\bm{X}\leftarrow\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{K}$ with Eq.(5) on $\bm{y}_{\text{p}}$ and $\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$
22:  end if
23:end for
Bidirectional Generation on Design-Score Samples via Derivative-Free Guidance
@@
In terms of score-to-design generation, the proposed method enables guided generation of preferred design-score samples through two mechanisms: (1) leveraging preferred scores $\bm{y}_{\text{p}}$ to guide the score component of the generated samples, ensuring alignment with desired performance; and (2) specifying a design constraint range $\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$ to guide the design component, ensuring design feasibility. This dual-guidance strategy generates samples that satisfy both design and score constraints without requiring an additional surrogate score model.
Concretely, during the reverse process, the guidance on the score component of the generated sample $\hat{\bm{x}}_{t}=(\bm{x}_{t},\bm{y}_{t})$ is formulated as the gradient of the MSE between $\bm{y}_{\text{p}}$ and the current score $\bm{y}_{t}$, i.e., $\nabla_{\bm{y}_{t}}\|\bm{y}_{\text{p}}-\bm{y}_{t}\|_{2}^{2}$. Similarly, the guidance on the design component is formulated as the gradient of the MSE from the current design $\bm{x}_{t}$ to the constraint range $\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}}]$, i.e., $\nabla_{\bm{x}_{t}}\|\bm{x}_{t}-\text{clip}(\bm{x}_{t},\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}})\|_{2}^{2}$, where $\text{clip}(\bm{x}_{t},\bm{x}_{\text{c}_{\text{min}}},\bm{x}_{\text{c}_{\text{max}}})$ ensures that $\bm{x}_{t}$ is projected onto the constraint range.
@@
where $\mathcal{F}(\bm{s}_{\bm{\theta}})\in(0,1]$, $j_{i}^{*}=\arg\min_{j}\|{\bm{x}}_{\text{uc}}^{(i)}-\bm{x}_{\text{train}}^{(j)}\|_{2}$ denotes the index of the nearest training data to ${\bm{x}}_{\text{uc}}^{(i)}$ in $\mathcal{D}=\{\hat{\bm{x}}^{(j)}_{\text{train}}=(\bm{x}^{(j)}_{\text{train}},\bm{y}^{(j)}_{\text{train}})\}_{j=1}^{N}$. A higher value of $\mathcal{F}(\bm{s}_{\bm{\theta}})$ indicates a stronger alignment between the generated and ground-truth scores. As shown in Figure 4, this metric exhibits a strong correlation with inference-time scaling performance based on the self-supervised reward mechanism. Based on this observation, we implement a conditional logic with $\mathcal{F}(\bm{s}_{\bm{\theta}})>\tau$ to activate the inference-time scaling scheme during guidance generation. In this way, we establish a general inference-time scaling framework for offline optimization by substituting external reward models in existing schemes with self-supervised rewards.
This framework is compatible with various methods, including the IS-based scaling [52] and the FKS-based scaling [53] methods. For example, we illustrate the complete process of ManGO based on IS with the self-supervised reward in Algorithm 1. Self-IS-based ManGO is designed to explore the denoising paths with higher rewards. Specifically, the algorithm first duplicates each noise point $\hat{\bm{x}}_{t}^{(i)}$ $J$ times. Each copy then undergoes an independent denoising step as defined in Eq. (5), where the inherent randomness in the reverse process—introduced through the Wiener noise term. This causes the denoising paths to diverge and results in divergent candidates $\{\hat{\bm{x}}_{t-1}^{(i,j)}\}_{j=1}^{J}$ at the next timestep. For each candidate, ManGO estimates its corresponding denoised state to compute self-supervised rewards $\mathcal{R}(\hat{\bm{x}}_{t-1}^{(i,j)})$ Eq.s (8) and (9). These rewards drive an importance sampling mechanism that prioritizes higher-reward paths, effectively steering the generative process toward more promising regions of the design-score manifold. This framework leverages the innate stochasticity of diffusion models to explore the noise space dynamically, eliminating the need for explicit external perturbations or reward models. Our experiments validate the integration of the IS and FKS approaches with self-supervised rewards.
Training and Inference Settings
@@ Acknowledgements
The ManGO’s main idea and training part were done during Tailin’s internship at Huawei Noah’s Ark Lab, and the inference part was done at HKUST. This work was supported in part by the Hong Kong Research Grants Council under the Areas of Excellence scheme grant AoE//E-601//22-R, in part by NSFC//RGC Collaborative Research Scheme grant CRS_\_HKUST603//22, in part by Guangzhou Municipal Science and Technology Project under Grant 2023A03J0011, Guangdong Provincial Key Laboratory of Integrated Communications, Sensing and Computation for Ubiquitous Internet of Things, and National Foreign Expert Project, Project Number G2022030026L. Some icons in Figure 1 are from the website https://www.cleanpng.com/.